# Teaching an Old Dog New Tricks: Verifiable FHE Using Commodity Hardware

### v1

@@ 4.1. FHE Applications Do Not Expose Secrets
No Secrets Exposed. In standard TEE platform, a central feature is to enforce the confidentiality of the execution environment. In FHE, all data is encrypted. As a result, the content of our enclave program (i.e., the state of ) is public. This means our application is not vulnerable to side channels, and we do not need to harden the execution environment against the usual threats.
No Long-Term Secret Storage. In usual TEE platforms, an important feature is the ability for an enclave program to seal secrets and recover them later. This property is usually tied to the identity (or the measurement) of the enclave program, and a secret can only be unsealed by a program that matches the correct measurement. This is, for instance, what is used in Bitlocker to only ever release the disk encryption key to a correctly booted system. TPMs and TEE platforms such as SGX all offer a seal operation. FHE evaluation does not require long-term secret storage, which significantly reduces the features required by the TEE.
No Need To Establish Trust. In usual remote attestation protocols, the remote client first needs to verify the TEE attestation to establish trust. Only once it trusts the TEE to have been correctly setup will it send its encrypted private input. This is because the inputs will then be decrypted inside of the TEE. In our case, inputs are never decrypted, which means that the remote client does not need to establish trust before sending the encrypted input. This makes it possible to amortize the cost of attestation by only signing the final transcript of program inputs and outputs (see section 5).
@@ 4.2. Eliminating Microarchitectural Side Channels Using A Physical TPM
Because our FHE-applications do not expose secrets, the only secret ever present on the server is the private signing key used in the attestation scheme. This is a much smaller attack surface than in usual TEE systems. We can take advantage of that opportunity to completely eliminate microarchitectural side channels by placing the signing key in a microarchitecturally isolated physical TPM. Because no secrets are left on the CPU or in main memory, this effectively makes our platform completely secure against microarchitectural side channels.
Side Channels Are Read-Only. One important element to keep in mind is that side channels are “read-only” gadgets and can only extract secrets from a victim program. Specifically, a side channel cannot modify the state of a victim’s program memory or change its control flow. These attacks are called fault injections and have not been as successful at attacking TEE platforms (see section 9.7).
@@
Physical TPMs Are Microarchitecturally Isolated. TPM are commonly implemented on a discrete chip or as firmware on a secure co-processor (we cannot secure software TPM running on the CPU). That means they are communicating with the main CPU using a hardware bus. They do not share cache nor any processor resources with the CPU. They use their own private resources, like private memory and registers, that are only accessible to the TPM. As a result, if the secret key is kept in a physical TPM, it is microarchitecturally isolated from the attacker and there is no possibility for the secret to be extracted through a microarchitectural side channel. Intrusive physical attacks might still be possible but are outside of our threat model (see section 9.7).
@@ 4.3. System Overview
Our system is composed of several hardware and software elements (see fig. 1). The security monitor is a small (12KLOC) trusted piece of code running at the hypervisor privilege level. It is in charge of enforcing isolation between the different security domains and integrity of the enclave programs. Isolation is enforced by hardware mechanisms (see section 8.2) and integrity is enforced through our attestation scheme (section 5). The security monitor does not manipulate any secret key material and delegates all cryptographic operations that require a key to the physical TPM. Integrity of the security monitor is enforced by the dynamic root-of-trust mechanism (section 8.1). Running on top of the security monitor, a large untrusted operating system is in charge of all resource allocations. It contains all the complex logic, for instance, for memory allocation or scheduling. It also provides many services for (enclave) applications such as networking, or IO.
Running an Enclave Program. To setup an trusted execution environment and securely run an enclave program, the untrusted OS interacts with the security monitor through vmcalls. On behalf of the OS, the security monitor will reclaim some resources (memory and cores) and allocate them to the enclave. Thanks to the isolation mechanisms under the control of the security monitor, the OS is now unable to modify the state of the enclave memory. The security monitor will also enforce some invariants (setting the page table correctly and zeroing regions where no content is loaded), ensure the correct binary is loaded, and attest the enclave initial state (see section 5.4). Once loaded, the enclave can be started. The remote client sends inputs over the network. They are received by the OS and placed in shared memory so that the enclave can access them. Inputs are systematically copied to enclave’s private memory before being used. Any output is also copied to shared memory and sent back to the client by the OS.
@@ 5.1. Existing Attestation Protocols
Attestation schemes are generally interactive protocols that involve several rounds of communication between a client and a server. First, before any sensitive data is sent, trust needs to be established between the execution environment running server side and the remote client. The client will send a challenge and receive an attestation report, endorsed by the manufacturer and attesting that a genuine piece of hardware has been correctly set up to instantiate a trusted execution environment. Once trust has been established, the enclave program and the remote client can perform key exchange and establish a trusted communication channel. They can then start communicating and running the expected application. Any message coming out of the enclave needs to be encrypted and attested using the private key generated inside of the enclave to prevent a malicious attacker from observing or tempering with the content of the messages. Similarly, every message from the remote client needs to be decrypted and its integrity verified.
@@ 5.3. Attested Transcript for FHE Applications
Trust does not need to be established before sending encrypted data. Because intermediate results are also encrypted, privacy cannot be compromised until a ciphertext is decrypted. That means integrity only needs to be checked just before decrypting an output (see section 6 for our security analysis). In other words, each message of the FHE application does not need to be individually attested. Instead, the integrity-protected security monitor can keep a transcript of all relevant data (i.e. all application inputs, intermediate results, and final output), and only perform a single TPM signature over that transcript. This simplifies the attestation scheme and drastically reduces the cost of using a physical TPM as remote attestation is now a simple fixed cost.
@@ 5.4. Argos Attestation Scheme
Measured Boot. Upon boot, the hardware DRoT is used to derive a measurement of the Argos security monitor (see section 8.1). This measurement is securely stored in the TPM registers (PCRs).
Enclave Setup. When a new enclave is initialized, the security monitor will allocate some of its private memory to the enclave transcript. During enclave set up, the security monitor measures the enclave binary and stores it in the transcript. In our specific case, the binary contains the FHE circuit, small public inputs (e.g., FHE parameters), alongside the logic to evaluate the circuit (the SEAL library in our case). It also contains the rest of the application logic. All the values of the relevant architectural registers that represent the initial state of the execution environment are also added to this measurement. For instance, the enclave initial program counter, page tables, the stack pointer and the addresses of shared memory use to communicate with the OS. The measurement of the binary and initial state unambiguously represent a function that takes inputs from shared memory and outputs results to it.
@@
Client Verification. Upon receiving the result(s) and the attested transcript, the client first verifies the attestation. The verification of the attestation goes as follows: First, the client verifies that the key used to sign the transcript belongs to a genuine TPM (see appendix A for more details). It then checks that the measurement for the security monitor contained in the transcript signature matches a reference measurement that is known to be correct (see section 9.7 on how such a reference measurement can be obtained). Finally, it checks that the binary references in the transcript corresponds to the correct FHE application and that it was evaluated on the input it provided. If all checks are satisfied, the verification is successful. In case of a verification failure, the server is behaving maliciously and the client aborts.
@@ 5.5. Remote Attestation as a Proof System
Additionally, our scheme satisfies these two properties:
Completeness. The system is complete if will always accept an honestly computed result i.e., a correctly formed transcript where .
@@
Security Analysis. Completeness of our attestation scheme reduces to the completeness of the underlying signature scheme used in the TPM. Soundness is more complex to establish, as it requires informally laying down all of our system security. First, as we explained earlier, an attacker (as described in our threat model) is not able to extract the private key from the TPM. Second, the security of our DRoT guarantees that the security monitor was correctly loaded and its integrity cannot be compromised. Third, our assumption of the TCB tells us that the security monitor is correctly implemented and bug-free. As a result, attested transcripts will only be generated for enclaves that are correctly set up, loaded, and executed. Our assumption on the underlying hardware isolation mechanisms (execution modes and nested page tables) guarantees that the integrity of the enclave code and data is maintained over the course of its lifetime. Finally, the functional correctness of the hardware guarantees that when an enclave is run that implements functionality on input , the resulting output . With all of this in place, the soundness of our attestation scheme reduces to the soundness of the TPM signature scheme.
@@ 6. Circuit-Level Verifiable FHE
In our client-server setup, all state-of-the-art FHE schemes (Chen et al., 2017a; Chillotti et al., 2020; Ducas and Micciancio, 2015) are insecure when considering a malicious server. Let us take a simple example where a client wants to outsource some sensitive computation to a server using FHE. Here, the server might be able to infer information regarding the decryption of a ciphertext by observing the client reaction. For instance, in FHE, if the noise in the ciphertext overflows a given threshold, client-side decryption might fail, which could lead to the client emitting a new request to the server. This is not a problem if assuming an honest-but-curious server. However, if the server is malicious, it can tweak the response to the client in arbitrary ways and observe its reaction. This is equivalent to providing our attacker with access to a (limited) decryption oracle that outputs “success” or “failure” for ciphertexts under the attacker’s control. This is already enough to mount key-recovery attacks (Guo et al., 2024; Chenal and Tang, 2015) and break all security for the FHE scheme (if the attacker recovers the private key, it can now decrypt the client’s private requests). In fact, this is much more problematic than a simple “correctness” issue where the client would just get a wrong output. In this simple outsourcing scenario, enforcing integrity at the FHE-circuit-level is enough, as it prevents the attacker from providing malformed ciphertexts to be decoded. When the client receives the response from the Argos server, it will first verify the attested transcript to ensure that the correct circuit was evaluated on the correct inputs. If verification fails, the client aborts, does not decrypt the ciphertext, and detects malicious behavior from the server. If and only if verification succeeds, the client may decrypt the ciphertext with the guarantee that it was correctly formed.
@@ 6.1. Definition
Correctness is defined identically to a semi-honest FHE scheme (see section 3.1). Security differs as our attacker is now malicious and can send arbitrary values. We also add two properties compared to the vanilla FHE-scheme.
@@ 6.2. Putting Everything Together
Let’s consider an FHE scheme (, , , ) and our remote attestation scheme (, , ). Argos’ construction is as follow:
@@ 7. Extending Argos to FHE-Based Applications
Circuit-level security sometimes differs from application-level security. For instance, attacks on correctness can escalate to attacks on privacy (section 3.3). Additionally, the security of our FHE schemes breaks if a malicious server is able to supply malformed ciphertexts for decryption section 3.3. This is potentially dangerous when, in many FHE-applications, the server itself can provide public or private inputs and act as a party in the computation. We explore two such applications, 1) private information retrieval, and 2) private set intersection. For each application, we discuss the challenges of deploying FHE-based schemes a real-world setting and show how Argos can be extended to enforce malicious security.
@@ 7.1. Malicious Private Information Retrieval
Private information retrieval (PIR) allows a client to retrieve an element from a public database hosted on a remote server, without the server learning which element was requested by the client. PIR has many applications such as public-key directory (Colombo et al., 2023), certificate transparency logs (Ryan, 2013; Henzinger et al., 2023b), private web search (Henzinger et al., 2023a) or private analytics (Green et al., 2016). PIR can be instantiated in a multi-server or single-server setting using a range of cryptographic primitives, but in this paper we focus on the state of the art for single-server PIR which are FHE-based schemes (Angel et al., 2018; Menon and Wu, 2022; Henzinger et al., 2023b; Menon and Wu, 2024).
@@
Defending Against Selective Failure Attacks. In these attacks, a malicious server chooses an entry of the database and replaces the entry with a value that it knows will cause a decryption failure. This is possible because in FHE, using an input that is not in the correct plaintext space (for instance, an integer that is too large) can lead to a malformed ciphertext which will not decrypt properly. Because our attacker might be able to observe the client reaction to decrypting the output (e.g., the client will send the request again), if decryption fails, the attacker can infer with high probability that the client queried the malformed entry.
@@
Authenticated PIR. So far, our malicious PIR scheme is only able to guarantee that the server performed the protocol using a well-formed database but not a specific one. Adding authenticity to a PIR scheme is a desirable property and might be essential for some real-world deployments. For instance, in the case of a transparency log or a public-key directory, it might be essential to guarantee the correct database was queried. We take inspiration from previous work and solve this problem by having the server initially commit to the database. This assumes a reference commitment is available for the client to fetch out-of-band.111This assumption is made by Authenticated PIR (Colombo et al., 2023) but recent works (Dietz and Tessaro, 2024; de Castro and Lee, 2024) have also shown how using a limited number of “validation queries”, the client can verify with high probability that the committed database is correct. Argos can easily be extended to support this functionality by adding a fresh commitment to the database (i.e. a hash or the root of a Merkle tree) to the attestation transcript. When receiving the response to its query, a client should first verify the attestation transcript, including that the database commitment matches the reference one. If the verification fails, it should abort. If (and only if) it succeeds, it should then decrypt the response and proceed. In our formalism, this is equivalent to extending Verify to take a reference commitment as an input, and tweaking soundness to ensure an adversary cannot make Verify accept a transcript that does not contain the correct commitment.

### v2

@@ 4.1. FHE Applications Do Not Expose Secrets
No Secrets Exposed. In a standard TEE platform, a central feature is to enforce the confidentiality of the execution environment. In FHE, all data is encrypted. As a result, the contents of our enclave program (i.e., the state of ) is public. This means our application is not vulnerable to side channels, and we do not need to harden the execution environment against the usual threats.
No Long-Term Secret Storage. In typical TEE platforms, an important feature is the ability for an enclave program to seal secrets and recover them later. This property is usually tied to the identity (or the measurement) of the enclave program, and a secret can only be unsealed by a program that matches the correct measurement. This is, for instance, what is used in Bitlocker to only ever release the disk encryption key to a correctly booted system. TPMs and TEE platforms such as SGX all offer a seal operation. FHE evaluation does not require long-term secret storage, which significantly reduces the features required by the TEE.
No Need To Pre-Establish Trust Before Sending Inputs. In usual remote attestation protocols, the remote client first needs to verify the TEE attestation to establish trust. Only once it trusts the TEE to have been correctly setup will it send its encrypted private input. This is because the inputs will then be decrypted inside of the TEE. In our case, inputs are never decrypted, which means that the remote client does not need to establish trust before sending the encrypted input. This makes it possible to amortize the cost of attestation by only signing the final transcript of program inputs and outputs (see Section 5).
@@ 4.2. Eliminating Microarchitectural Side Channels Using A Physical TPM
Because our FHE-applications do not expose secrets, the only secret ever present on the server is the private signing key used in the attestation scheme. This is a much smaller attack surface than in usual TEE systems. We can take advantage of that opportunity to completely eliminate microarchitectural side channels by placing the signing key in a microarchitecturally-isolated physical TPM. Because no secrets are left on the CPU or in main memory, this effectively makes our platform completely secure against microarchitectural side channels.
Side Channels Are Read-Only. One important element to keep in mind is that side channels are “read-only” gadgets and can only extract secrets from a victim program. Specifically, a side channel cannot modify the state of a victim’s program memory or change its control flow.
@@
Physical TPMs Are Microarchitecturally Isolated. TPMs are commonly implemented on a discrete chip or as firmware on a secure co-processor (we cannot secure software TPM running on the CPU). They communicate with the main CPU using a hardware bus. They do not share cache, nor any processor resources, with the CPU. They use their own private resources, like private memory and registers, that are only accessible to the TPM. As a result, if the secret key is kept in a physical TPM, it is microarchitecturally-isolated from the attacker and there is no possibility for the secret to be extracted through a CPU microarchitectural side channel. In addition, the microarchitecture of TPMs and secure co-processors is intentionally kept simple, avoiding features like caches or out-of-order execution. This prevents the existence of indirect microarchitectural side channels such as NetSpectre (Schwarz et al., 2019). Intrusive physical attacks might still be possible but are outside of our threat model (see Section 9.7).
@@ 4.3. System Overview
Our system is composed of several hardware and software elements (see Figure 1). The security monitor is a small (18 KLOC) trusted piece of code running at the hypervisor privilege level. It is in charge of enforcing isolation between the different security domains and integrity of the enclave programs. Isolation is enforced by hardware mechanisms (see Section 8.2) and integrity is enforced through our attestation scheme (Section 5). The security monitor does not manipulate any secret key material and delegates all cryptographic operations that require a key to the physical TPM. Integrity of the security monitor is enforced by the dynamic root-of-trust mechanism (Section 8.1). Running on top of the security monitor, a large untrusted operating system is in charge of all resource allocations. It contains all the complex logic, for instance, for memory allocation or scheduling. It also provides many services for (enclave) applications such as networking, or I/O.
Running an Enclave Program. To setup a trusted execution environment and securely run an enclave program, the untrusted OS interacts with the security monitor through vmcalls to the hypervisor. On behalf of the OS, the security monitor will reclaim some resources (memory and cores) and allocate them to the enclave. Thanks to the isolation mechanisms under the control of the security monitor, the OS is now unable to modify the state of the enclave memory. The security monitor will also enforce some invariants (setting the page table correctly and zeroing regions where no content is loaded), ensure the correct binary is loaded, and attest the enclave initial state (see Section 5.4). Once loaded, the enclave can be started. The remote client sends inputs over the network. They are received by the OS and placed in shared memory so that the enclave can access them. Inputs are systematically copied to enclave’s private memory before being used. Any output is also copied to shared memory and sent back to the client by the OS.
@@ 5.1. Existing Attestation Protocols
Attestation schemes are generally interactive protocols that involve several rounds of communication between a client and a server. First, before any sensitive data is sent, trust needs to be established between the execution environment running on the server side and the remote client. The client will send a challenge and receive an attestation report, endorsed by the manufacturer and attesting that a genuine piece of hardware has been correctly set up to instantiate a trusted execution environment. Once trust has been established, the enclave program and the remote client can perform key exchange and establish a trusted communication channel. They can then start communicating and running the expected application. Any message coming out of the enclave needs to be encrypted and attested using the private key generated inside of the enclave to prevent a malicious attacker from observing or tampering with the content of the messages. Similarly, every message from the remote client needs to be decrypted and its integrity verified.
@@ 5.3. Attested Transcript for FHE Applications
Trust does not need to be established before sending encrypted data. Because intermediate results are also encrypted, privacy cannot be compromised until a ciphertext is decrypted. That means integrity only needs to be checked just before decrypting an output (see Section 6 for our security analysis). In other words, each message of the FHE application does not need to be individually attested. Instead, the integrity-protected security monitor can keep a transcript of all relevant data (i.e., all application inputs, intermediate results, and final output), and, only perform a single TPM signature over that transcript. This simplifies the attestation scheme and drastically reduces the cost of using a physical TPM as remote attestation is now a simple fixed cost (see Figure 2).
@@ 5.4. Argos Attestation Scheme
Measured Boot. Upon boot, the hardware dynamic root of trust (DRoT) is used to derive a measurement of the Argos security monitor (see Section 8.1). This measurement is securely stored in the TPM registers (PCRs).
Enclave Setup. When a new enclave is initialized, the security monitor will allocate some of its private memory to the enclave transcript. During enclave setup, the security monitor measures the enclave binary and stores it in the transcript. In our specific case, the binary contains the FHE circuit, small public inputs (e.g., FHE parameters), alongside the logic to evaluate the circuit (the SEAL library in our case). It also contains the rest of the application logic. All the values of the relevant architectural registers that represent the initial state of the execution environment are also added to this measurement. For instance, the enclave initial program counter, page tables, the stack pointer and the addresses of shared memory used to communicate with the OS. The measurement of the binary and initial state unambiguously represent a function that takes inputs from shared memory and outputs results to it.
@@
Client Verification. Upon receiving the result(s) and the attested transcript, the client first verifies the attestation. The verification of the attestation goes as follows: First, the client verifies that the key used to sign the transcript belongs to a genuine TPM (see Appendix A for more details). It then checks that the measurement for the security monitor contained in the transcript signature matches a reference measurement that is known to be correct (see Section 9.7 on how such a reference measurement can be obtained). Finally, it checks that the binary reference in the transcript corresponds to the correct FHE application and that it was evaluated on the input it provided. If all checks are satisfied, the verification is successful. In case of a verification failure, the server is behaving maliciously and the client aborts.
@@ 5.5. Remote Attestation as a Proof System
Additionally, our scheme satisfies these two properties (see Appendix C.2 for formal definitions):
Completeness. The system is complete if will always accept a honestly computed result, i.e., a correctly formed transcript where .
@@
Security Analysis. Completeness of our attestation scheme reduces to the correctness of the underlying signature scheme used in the TPM. Soundness is more complex to establish, as it requires informally laying down all of our system security. First, as we explained earlier, an attacker (as described in our threat model) is not able to extract the private key from the TPM. Second, the security of our DRoT guarantees that the security monitor was correctly loaded and its integrity cannot be compromised. Third, our assumption of the TCB tells us that the security monitor is correctly implemented and bug-free. As a result, attested transcripts will only be generated for enclaves that are correctly set up, loaded, and executed. Our assumption on the underlying hardware isolation mechanisms (execution modes and nested page tables) guarantees that the integrity of the enclave code and data is maintained over the course of its lifetime. Finally, the functional correctness of the hardware guarantees that when an enclave is run that implements functionality on input , the resulting output . With all of this in place, the soundness of our attestation scheme reduces to the soundness of the TPM signature scheme.
@@ 6. Circuit-Level Verifiable FHE
In our client-server setup, all state-of-the-art FHE schemes (Chen et al., 2017a; Chillotti et al., 2020; Ducas and Micciancio, 2015) are insecure when considering a malicious server. Let us take a simple example where a client wants to outsource some sensitive computation to a server using FHE. Here, the server might be able to infer information regarding the decryption of a ciphertext by observing the client reaction. For instance, in FHE, if the noise in the ciphertext overflows a given threshold, client-side decryption might fail, which could lead to the client emitting a new request to the server. This is not a problem if assuming an honest-but-curious server. However, if the server is malicious, it can tweak the response to the client in arbitrary ways and observe its reaction. This is equivalent to providing our attacker with access to a (limited) decryption oracle that outputs “success” or “failure” for ciphertexts under the attacker’s control. This is already enough to mount key-recovery attacks (Guo et al., 2024; Chenal and Tang, 2015) and break security of the FHE scheme (if the attacker recovers the private key, it can now decrypt the client’s private requests). In fact, this is much more problematic than a simple “correctness” issue where the client would just get a wrong output. In this simple outsourcing scenario, enforcing integrity at the FHE-circuit-level is enough, as it prevents the attacker from providing malformed ciphertexts to be decoded. When the client receives the response from the Argos server, it will first verify the attested transcript to ensure that the correct circuit was evaluated on the correct inputs. If verification fails, the client aborts, does not decrypt the ciphertext, and detects malicious behavior from the server. If and only if verification succeeds, the client may decrypt the ciphertext with the guarantee that it was correctly formed.
@@ 6.1. Definition
Formal properties can be found in Appendix C.3. Correctness is defined identically to a semi-honest FHE scheme (see Section 3.1). Security differs as our attacker is now malicious and can send arbitrary values. We also add two properties compared to the vanilla FHE-scheme.
@@ 6.2. Putting Everything Together
Let’s consider an FHE scheme (, , , ) and our remote attestation scheme (, , ). Argos’ construction is as follows:
@@ 7. Extending Argos to FHE-Based Applications
Circuit-level security sometimes differs from application-level security. For instance, attacks on correctness can escalate to attacks on privacy (Section 3.3). Additionally, the security of our FHE schemes breaks if a malicious server is able to supply malformed ciphertexts for decryption (Section 3.3). This is potentially dangerous when, in many FHE-applications, the server itself can provide public or private inputs and act as a party in the computation. We explore two such applications, 1) private information retrieval, and 2) private set intersection. For each application, we discuss the challenges in deploying FHE-based schemes in a real-world setting and show how Argos can be extended to enforce malicious security.
@@ 7.1. Malicious Private Information Retrieval
Private information retrieval (PIR) allows a client to retrieve an element from a public database hosted on a remote server, without the server learning which element was requested by the client. PIR has many applications such as public-key directory (Colombo et al., 2023), certificate transparency logs (Ryan, 2013; Henzinger et al., 2023b), private web search (Henzinger et al., 2023a) or private analytics (Green et al., 2016). PIR can be instantiated in a multi-server or single-server setting using a range of cryptographic primitives, but in this paper we focus on the state-of-the-art for single-server PIR which are FHE-based schemes (Angel et al., 2018; Menon and Wu, 2022; Henzinger et al., 2023b; Menon and Wu, 2024).
@@
Defending Against Selective Failure Attacks. In these attacks, a malicious server chooses an entry in the database and replaces the entry with a value that it knows will cause a decryption failure. This is possible because in FHE, using an input that is not in the correct plaintext space (for instance, an integer that is too large) can lead to a malformed ciphertext, which will not decrypt properly. Because our attacker might be able to observe the client reaction to decrypting the output (e.g., the client will send the request again), if decryption fails, the attacker can infer with high probability that the client queried the malformed entry.
@@
Authenticated PIR. So far, our malicious PIR scheme is only able to guarantee that the server performed the protocol using a well-formed database, but not a specific one. Adding authenticity to a PIR scheme is a desirable property and might be essential for some real-world deployments. For instance, in the case of a transparency log or a public-key directory, it might be essential to guarantee the correct database was queried. We take inspiration from previous work and solve this problem by having the server initially commit to the database. This assumes a reference commitment is available for the client to fetch out-of-band.222This assumption is made by Authenticated PIR (Colombo et al., 2023) but recent works (Dietz and Tessaro, 2024; de Castro and Lee, 2024) have also shown how using a limited number of “validation queries”, the client can verify with high probability that the committed database is correct. Argos can easily be extended to support this functionality by adding a fresh commitment to the database (i.e., a hash or the root of a Merkle tree) to the attestation transcript. When receiving the response to its query, a client should first verify the attestation transcript, including that the database commitment matches the reference one. If the verification fails, it should abort. If (and only if) it succeeds, it should then decrypt the response and proceed. In our formalism, this is equivalent to extending Verify to take a reference commitment as an input, and tweaking soundness to ensure an adversary cannot make Verify accept a transcript that does not contain the correct commitment. A detailed security argument can be found in Appendix E.